# SpectrAx: Spectral Search of Axion-Like Particles Using Multi-Band Observations of Galaxy Clusters from SKA, SO, CMB-S4 and eROSITA

### v1

@@ Abstract
The existence of axions or Axion-Like Particles (ALPs) has been predicted by various Beyond Standard Model (BSM) theories, and the proposed photon-ALP interaction is one of the ways to probe them. Such an interaction will lead to photon-ALP resonant conversion in galaxy clusters, resulting in a polarized spectral distortion in the CMB along the cluster line of sight. The estimation of this signal from galaxy clusters requires an estimation of the electron density and magnetic field in galaxy clusters. We have developed a new Bayesian framework SpectrAx that can use observations from different electromagnetic bands such as radio, CMB, optical, and X-ray to infer the astrophysical properties of a galaxy cluster, such as cluster its redshift, electron density and magnetic field, along with the BSM physics such as ALPs. By using the simulated data for upcoming CMB surveys such as Simons Observatory (SO) and CMB-S4 in combination with Square Kilometer Array (SKA) and extended ROentgen Survey with an Imaging Telescope Array (eROSITA) we demonstrate the capability in accurately inferring the ALPs coupling strength along with the radial profile of electron density and magnetic field from galaxy clusters. The application of this framework to the data from future surveys by combining SKA+SO+eROSITA and SKA+CMB-S4+eROSITA will make it possible for the first time to explore both astrophysics and BSM physics from low-redshift galaxy clusters using a multi-band approach.
@@ 1 Introduction
The $\Lambda$CDM universe consists of various components such as radiation, matter, and dark energy, with all of these affecting the evolution of the universe during different stages [1, 2, 3]. The matter component is composed mostly of dark matter ($\sim 85\%$), followed by baryons ($\sim 15\%$), with the rest being non-relativistic neutrinos. Dark matter halos are the largest gravitationally bound structures in the universe. These halos are hosts to galaxy clusters, which are the largest observable gravitationally bound structures. Thus, it is these clusters that can help us connect astrophysics to cosmology. These clusters can be observed due to the multiple phenomena giving rise to different kinds of emissions, corresponding to various baryon and baryon-photon interactions [4, 5, 6, 7].
@@
The motivation behind developing the multi-band framework is highlighted in Sec. 2. This is followed by a review of the various emissions and the photon-ALP resonant conversion phenomenon relevant for SpectrAx in Sec. 3. The components of SpectrAx and their working is explained in Sec. 4. The results on the coupling constant, obtained from mock data using SpectrAx are showcased in Sec. 5. The multi-band framework and its significance in the upcoming era are summarized in Sec. 6. We have mostly used natural units ($\hbar=1,c=1,k_{B}=1$), until explicitly mentioned and the cosmological parameters from Planck 2015 results [47].
@@ 2 Motivation
The galaxy clusters are the largest observable gravitationally bound structures in the universe, and act as a bridge between astrophysics and cosmology. The intracluster medium (ICM) and the galaxies are sites for many electromagnetic phenomena occurring in clusters [48]. These lead to emissions that cover a broad range of the EM spectrum. Fig. 1 shows the emission from a galaxy cluster at various frequencies (low to high). The cluster emits mainly through synchrotron radiation on the left at radio frequencies, and through thermal X-ray emission on the right at high frequencies. The emission at the center is in the presence of CMB, if ALPs exist in nature, which will lead to an ALP distortion in the CMB in the cluster region.
Figure 1: The emission from a cluster at different frequencies. The cluster emits mainly through synchrotron radiation on the left at radio frequencies, and through thermal X-ray emission on the right at high frequencies. The emission at the center is in the presence of CMB, if ALPs exist in nature, which will lead to an ALP distortion in the CMB in the cluster region.
@@
The framework SpectrAx, developed in this work brings the astrophysics of clusters and the cosmological search for dark matter (axions being one of the candidates) under one umbrella, using multi-band observations of galaxy clusters. Since various emissions from a cluster depend on different properties of a cluster, all the cluster attributes cannot be inferred from a single frequency band, dominated by just a particular kind of emission. Multi-frequency analysis enables a study of different emissions from clusters. The ALP signal generated in galaxy clusters depends on the magnetic field and electron density profiles, and the redshift of the cluster, which can be inferred across different frequency surveys. The optical observations can provide the redshift information, while the thermal X-ray emission can probe the electron density and temperature profiles of a cluster. The radio emissions from synchrotron will provide constraints on the transverse magnetic field profile. Using CMB observations (contaminated with foregrounds and instrument noise) of cluster regions, and cleaning them, the ALP signal can be probed to obtain constraints on the photon-ALP coupling constant. The inferred properties from different bands can then be combined under the hierarchical Bayesian framework to get constraints on the coupling constants of ALPs (see Fig. 2).
Observational probe Bandwidth Properties ascertained
Radio 700 MHz to 15 GHz Magnetic field
@@
Optical 400 to 700 THz Cluster redshift
@@ 3.1 Magnetic field from radio observations
For our analysis, we consider a radial profile for the magnetic field in galaxy clusters that decays in the outer regions of the cluster. We calculate the line of sight integrated synchrotron intensity (using Equ.(3.3)) at scales corresponding to the pixel resolution and smooth with the beam size of the instrument. This is done at different frequency bands to obtain a multi-frequency-based magnetic field inference. The instrument noise is added corresponding to the frequency bands to generate a radio synchrotron emission map for galaxy clusters. The intensity values in the cluster region will provide us with the data vector for our Bayesian analysis. This will allow us to perform a radial tomographic reconstruction of the magnetic field profile using radio observations. This analysis will enable a better understanding of magnetic fields in galaxy clusters and their coherence scales. This would be key in being able to estimate the 3-dimensional profile for non-spherical clusters. The inference of a three-dimensional magnetic field taking into account the coherence, as well as the turbulence in the magnetic field using synchrotron, as well as the Faraday rotation measurements will be included in the SpectrAx framework in the future. In this analysis, we focus on the inference of transverse magnetic field using synchrotron emission.
@@ 3.2 Electron density and Temperature from X-rays
For low energy photons, the Thomson scattering takes place with optical depth $\tau\sim n_{e}\sigma_{T}R_{\mathrm{eff}}\sim 10^{-2}$ to $10^{-3}$ [79]. The SZ is a redshift-independent effect and can be used to identify clusters even at very high redshifts as it causes fractional change in CMB brightness. The SZ effect leads to an increase in brightness at higher frequencies, while a decrease at lower frequencies. The spectral dependence of the temperature fluctuations from SZ effect is given as (where $x\equiv h\nu/k_{B}T_{\mathrm{cmb}}$ ) [78, 80, 81, 13]:
@@
The varying dependence of X-rays and SZ effect on the electron density and temperature profiles can break the degeneracy in the inferred electron density and temperature amplitudes [15, 82, 83, 77]. Mass measurements from lensing can also improve the constraints on these profiles. We have not considered this in the current analysis, and can be included in SpectaAx framework in the future.
@@ 3.3 Cross-matching with the optical survey
The morphology and ellipticity of clusters can also be inferred from optical observations. Optical surveys are also used to provide us with velocity dispersion measurements which gives us constraints on the masses of clusters. The mass of a cluster can also be found using observations from optical, X-rays, and lensing [90, 74, 27]. The mass cross-matching will provide constraints on the mass of a cluster. This is currently not considered in the analysis.
3.4 Axion detection using CMB as a backlight
@@ 4 The SpectrAx pipeline
Notation Description Prior range
@@
A Gaussian likelihood has been considered with uniform priors on all parameters within the allowed parameter range in the case of radio and X-ray inference. During the MCMC parameter estimation, the burn in part of the chain is discarded and thinning is performed to minimize the correlation between steps.
To demonstrate the working of SpectrAx we perform a mock analysis for 10 mock resolved galaxy clusters, for which we can infer the magnetic field, electron density and redshift from radio, X-rays and optical observations. We consider spherical symmetry for these clusters and assume that the coherence scale of the magnetic field is resolvable, with the high-resolution beam of upcoming detectors. We will find constraints on the ALP coupling constant for ALPs of mass $10^{-14}$ eV (with 15% tolerance) from these clusters for a combination of SKA, eROSITA, and SO (or CMB-S4) configuration.
@@ 4.2 Framework for inference of the cluster magnetic field using radio observations
We use the dependence of the synchrotron emission on the magnetic field and spectral index (see Equ.(3.3)). Also, a proportionality factor is introduced, which we denote as $C$ and is varied for the clusters to generate signals of about $\sim 0.1$ to 2 $\mu\mathrm{Jy/arcsec^{2}}$ [65]. The profile parameters ($s$ and $B_{0}$ in this case), proportionality factor ($C$), and spectral index ($p$) are varied within the allowed range for galaxy clusters to generate random profiles for different clusters. We make the synchrotron maps by calculating the synchrotron emission integrated along the line of sight through the cluster for each of these clusters, integrated across the required frequency bands.
@@
We have assumed that the magnetic field varies over the angular scale, which is resolvable with the beam resolution of SKA. We have considered a radial profile, but in reality, the clusters will have a different geometry and the profiles will not be radial. This would require complex models with a 3-dimensional variation of the magnetic field. The inference of the magnetic field using synchrotron polarization measurements will provide better constraints on the magnetic field. Using the Faraday rotation $RM$ measurements, the magnetic field profile along the line of sight can also be constrained, which will be able to provide a 3-dimensional inference of the magnetic field direction and magnitude within the cluster. This can be included in an upgraded version of the SpectrAx code.
@@ 4.3 Framework for inference of the cluster electron density using X-rays
The estimation of 6 parameters ($n0$, $r_{c1}$, $\beta_{1}$, $T_{e}$, $\Gamma$ and $A$) using X-ray emission is shown is Fig. 6. The average time for the run is $\approx 35$ core-hours per 1000 pixels. The electron density and temperature profiles involve a large number of parameters and some of them portray degeneracies. The constraints depend on the strength of the signal as well as the number of pixels that the cluster occupies. The 95% confidence region for the inference of electron density for the cluster is shown in Fig. 7.
@@
We have assumed that the electron density varies over the angular scale, which is resolvable with the beam resolution of eROSITA. We have considered a radial profile, but in reality the clusters will have a different geometry and the profiles will not be radial. This would require complex models with 3-dimensional variation of electron density. SpectrAx works efficiently in estimating a large number of parameters, given computer resources and resolutions are available. The constrained parameters will be marginalized along with the magnetic field parameters and redshift to obtain the ALP coupling constant under the hierarchical Bayesian framework in Sec. 4.6.
4.4 Axion signal in CMB map

### v3

@@ Abstract
The existence of axions or Axion-Like Particles (ALPs) has been predicted by various Beyond Standard Model (BSM) theories, and the proposed photon-ALP interaction is one of the ways to probe them. Such an interaction will lead to photon-ALP resonant conversion in galaxy clusters, resulting in a polarized spectral distortion in the CMB along the cluster line of sight. The estimation of this signal from galaxy clusters requires an estimation of their electron density and magnetic field profiles, as well as their redshifts. We have developed a new Bayesian framework SpectrAx that can use observations from different electromagnetic bands such as radio, CMB, optical, and X-ray to infer the astrophysical properties of a galaxy cluster, such as cluster its redshift, electron density and magnetic field, along with the BSM physics such as ALPs. We use simulated redshifts in our analysis, but that can be obtained by cross-matching with optical surveys having overlapping sky regions with the galaxy clusters. Also, we use radial profiles that are motivated from observations of galaxy clusters at low redshifts. By using the simulated data corresponding to the ALP mass of $10^{-14}$ eV for upcoming CMB surveys such as Simons Observatory (SO) and CMB-S4 in combination with Square Kilometer Array (SKA) and extended ROentgen Survey with an Imaging Telescope Array (eROSITA) we demonstrate the capability in accurately inferring the ALPs coupling strength along with the radial profile of electron density and magnetic field from galaxy clusters. The application of this framework to the data from future surveys by combining SKA+SO+eROSITA and SKA+CMB-S4+eROSITA will make it possible for the first time to explore both astrophysics and BSM physics from low-redshift galaxy clusters using a multi-band approach.
@@ 1 Introduction
The $\Lambda$CDM universe consists of various components in radiation, matter, and dark energy, with all of these affecting the evolution of the universe during different stages [1, 2, 3]. The matter component is composed mostly of dark matter ($\sim 85\%$), followed by baryons ($\sim 15\%$), with the rest being non-relativistic neutrinos. Dark matter halos are the largest gravitationally bound structures in the universe. These halos are hosts to galaxy clusters, which are the largest observable gravitationally bound structures. Thus, it is these clusters that can help us connect astrophysics to cosmology. These clusters can be observed due to the multiple phenomena giving rise to different kinds of emissions, corresponding to various baryon and baryon-photon interactions [4, 5, 6, 7].
@@
The motivation behind developing the multi-band framework is highlighted in Sec. 2. This is followed by a review of the various emissions and the photon-ALP resonant conversion phenomenon relevant for SpectrAx in Sec. 3. The components of SpectrAx and their working is explained in Sec. 4. The results on the coupling constant, obtained from mock data using SpectrAx are showcased in Sec. 5. The multi-band framework and its significance in the upcoming era are summarized in Sec. 6. We have mostly used natural units ($\hbar=1,c=1,k_{B}=1$), until explicitly mentioned and the cosmological parameters from Planck 2015 results [47] (Planck TT, TE, EE + low P + Lensing data) for this analysis.
@@ 2 Motivation
The galaxy clusters are the largest observable gravitationally bound structures in the universe, and act as a bridge between astrophysics and cosmology. The intracluster medium (ICM) and the galaxies are sites for many electromagnetic phenomena occurring in clusters [48]. These lead to emissions that cover a broad range of the EM spectrum. Fig. 1 shows the emission from a galaxy cluster at various frequencies (low to high). The cluster emits mainly through synchrotron radiation on the left at radio frequencies, and through thermal X-ray emission on the right at high frequencies. The emission at the center shows the cluster at the microwave frequency of 145 GHz, if ALPs exist in nature. This will lead to an ALP distortion in the CMB in the cluster region. The angular scale is the same for all the three frequencies.
Figure 1: The emission from a cluster at different frequencies. The cluster emits mainly through synchrotron radiation on the left at radio frequencies, and through thermal X-ray emission on the right at high frequencies. The emission at the center shows the cluster at the microwave frequency of 145 GHz, if ALPs exist in nature. This will lead to an ALP distortion in the CMB in the cluster region. The black dashed curve encloses the resolved cluster line of sight. The scale is same for all the three subfigures.
@@
The framework SpectrAx, developed in this work brings the astrophysics of clusters and the cosmological search for axions or ALPs (one of the dark matter candidates), under one umbrella, using multi-band observations of galaxy clusters. Since various emissions from a cluster depend on different properties of a cluster, all the cluster attributes cannot be inferred from a single frequency band, dominated by just a particular kind of emission. Multi-frequency analysis enables a study of different emissions from clusters. The ALP signal generated in galaxy clusters depends on the magnetic field and electron density profiles, and the redshift of the cluster, which can be inferred across different frequency surveys (see Table 1). The optical observations can provide the redshift information, while the thermal X-ray emission can probe the electron density and temperature profiles of a cluster. The radio emissions from synchrotron will provide constraints on the transverse magnetic field profile. Using CMB observations (contaminated with foregrounds and instrument noise) of cluster regions, and cleaning them, the ALP signal can be probed to obtain constraints on the photon-ALP coupling constant. The inferred properties from different bands can then be combined under the hierarchical Bayesian framework to get constraints on the coupling constants of ALPs (see Fig. 2).
Probes Bandwidth Properties ascertained Considered in SpectrAx
Radio 0.7 to 15 GHz Magnetic field
Smooth radial profiles,
Synchrotron intensity
@@
Smooth radial profiles,
Foregrounds,
Variable ALP mass & coupling
Optical 400 to 700 THz Cluster redshift Simulated Redshifts
@@ 3.1 Magnetic field from radio observations
For our analysis, we consider a radial profile for the magnetic field in galaxy clusters that decays in the outer regions of the cluster. The magnetic field profile considered is motivated from observations of low redshift galaxy clusters (see Equ. (4.4)), but SpectrAx will be able to employ complicated magnetic field models as well in its code and estimate the parameters. The addition of nuisance parameters to the profiles can take into account the uncertainty in their modeling. This would make the analysis robust, but with increased computational cost. The addition of turbulence and incoherence would further add to this uncertainty, but these effects will be considered in a future analysis.
We have used different parameter values for different clusters within the allowed range for galaxy clusters. We calculate the line of sight integrated synchrotron intensity (using Equ.(3.3)) at scales corresponding to the pixel resolution and smooth with the beam size of the instrument. Depolarization occurs if the beam size is greater than the magnetic field coherence scale. We assume that is not the case with upcoming high resolution experiments like SKA. This is done at different frequency bands to obtain a multi-frequency-based magnetic field inference. The instrument noise is added corresponding to the frequency bands to generate a radio synchrotron emission map for galaxy clusters. The intensity values in the cluster region will provide us with the data vector for our Bayesian analysis. This will allow us to perform a radial tomographic reconstruction of the magnetic field profile using radio observations. This analysis will enable a better understanding of magnetic fields in galaxy clusters and their coherence scales. This would be key in being able to estimate the 3-dimensional profile for non-spherical clusters. The inference of a three-dimensional magnetic field taking into account the coherence, as well as the turbulence in the magnetic field using synchrotron, as well as the Faraday rotation measurements will be included in the SpectrAx framework in the future. In this analysis, we focus on the inference of radial transverse magnetic field using synchrotron emission, while the effects of turbulence and incoherence will be considered in a future work.
@@ 3.2 Electron density and Temperature from X-rays
For low energy photons, the Thomson scattering takes place with optical depth $\tau\sim n_{e}\sigma_{T}R_{\mathrm{eff}}\sim 10^{-2}$ to $10^{-3}$ [79]. The SZ is a redshift-independent effect and can be used to identify clusters even at very high redshifts as it causes a fractional change in CMB brightness. The SZ effect leads to an increase in brightness at higher frequencies, while a decrease at lower frequencies. The spectral dependence of the temperature fluctuations from SZ effect is given as (where $x\equiv h\nu/k_{B}T_{\mathrm{cmb}}$ ) [78, 80, 81, 13]:
@@
The varying dependence of X-rays and SZ effect on the electron density and temperature profiles can break the degeneracy in the inferred electron density and temperature amplitudes [15, 82, 83, 77]. Mass measurements from lensing can also improve the constraints on these profiles. We have not considered the SZ and mass measurements in the current analysis, and can be included in SpectaAx framework in the future.
@@ 3.3 Cross-matching with the optical survey
The morphology and ellipticity of clusters can also be inferred from optical observations. Optical surveys are also used to provide us with velocity dispersion measurements which gives us constraints on the masses of clusters. The mass of a cluster can also be found using observations from optical, X-rays, and lensing [90, 74, 27]. The mass cross-matching will provide constraints on the mass of a cluster, but is currently not considered in the analysis. This can be included in the next version of this pipeline.
3.4 Axion-Like Particles
Axions or ALPs (axion-like particles) are hypothetical particles, predicted by various beyond Standard model theories. Axions belong to the class of pseudo Nambu-Goldstone bosons (PNGBs) which possess mass. These were first proposed to be produced as solutions to the strong CP problem [91]. The global Peccei-Quinn chiral U(1) symmetry breaking (which is not exact at low energies) by the vacuum expectation value of a scalar field leads to the production of axions which possess a small mass, which also makes them one of the dark matter candidates [30, 31, 32, 33, 34, 35, 36, 37, 92].
Similar to the case of neutrino oscillations, axion-photon oscillations are possible if axions can weakly interact with photons [93, 94, 95, 96, 97, 98, 99, 100, 101, 102]. Since axions require high energies for production in particle-based experiments, detecting them using the weak coupling they may be having with photons makes for a suitable alternative [41, 38, 42, 103]. Galaxy clusters provide a way of probing ALP masses of a certain range ($\sim 10^{-15}-10^{-11}$ eV), but there is a wide range of ALP masses that can be probed using other astrophysical systems as well, such as neutrino halos, galactic halos, etc., [104, 105, 106, 107].
The strength of the interaction between a photon and ALP is given by a photon-ALP coupling ($g_{a\gamma}$), the variation of which with axion mass depends on the model. But in the case of ALPs, this variation is not well constrained. The bounds from COBE FIRAS on the product of magnetic field and coupling constant are at $\mathrm{g_{a\gamma}B<10^{-13}\,GeV^{-1}\,nG}$ [108]. The 95% confidence interval (C.I.) on the coupling constant is at $\mathrm{\leq 6.6\times 10^{-11}\,GeV^{-1}}$ [109] from CAST. The RMS fluctuation bounds on non-resonant ALP distortions using Planck data is $<18.5\times 10^{-6}$ [44]. Additionally, constraints on the effective number of neutrino species (denoted by $N_{\mathrm{eff}}$) from SO and CMB-S4 can also constrain the couplings of axions or other BSM particles using CMB anisotropies [110, 111, 112].
3.5 ALP detection using CMB as a backlight
@@ 4 The SpectrAx pipeline
Notation Description Uniform Prior range
@@
$p$ Synchrotron spectral index $93<p<$ 9
$C$ Synchrotron emission proportionality constant $15501450<C<$ 1550
@@
A Gaussian likelihood has been considered with uniform priors on all parameters within the allowed parameter range in the case of radio and X-ray inference. During the MCMC parameter estimation, the burn in part of the chain is discarded and thinning is performed to minimize the correlation between steps. The convergence for one of the clusters can be checked in Appendix A.
To demonstrate the working of SpectrAx we perform a mock analysis for 10 mock resolved galaxy clusters, for which we can infer the magnetic field, electron density and redshift from radio, X-rays and optical observations. We consider spherical symmetry for these clusters and assume that the coherence scale of the magnetic field is resolvable, with the high-resolution beam of upcoming detectors. We will find constraints on the ALP coupling constant for ALPs of mass $10^{-14}$ eV (with 15% tolerance about this mean to set the limits on the conversion length scale as the electron density profile decides the locations where different mass ALPs are formed, as per the resonant condition) from these clusters for a combination of SKA, eROSITA, and SO (or CMB-S4) configuration. Generally, the lower the ALP mass, the stronger is the expected ALP signal, hence better are the constraints.
@@ 4.2 Framework for inference of the cluster magnetic field using radio observations
We use the dependence of the synchrotron emission on the magnetic field and spectral index (see Equ.(3.3)). Also, a proportionality factor is introduced, which we denote as $C$ and is varied for the clusters to generate signals of about $\sim 0.1$ to 2 $\mu\mathrm{Jy/arcsec^{2}}$ [65]. The profile parameters ($s$ and $B_{0}$ in this case), proportionality factor ($C$), and spectral index ($p$) are assigned uniform priors (see Table 2) and varied within the allowed range for galaxy clusters to generate random profiles for different clusters. We make the synchrotron maps by calculating the synchrotron emission integrated along the line of sight through the cluster for each of these clusters, integrated across the required frequency bands.
@@
We have assumed that the magnetic field varies over the angular scale, which is resolvable with the beam resolution of SKA. We have considered a radial profile, but in reality, the clusters will have a different geometry and the profiles will not be radial. This would require complex models with a 3-dimensional variation of the magnetic field. The stochasticity and turbulence in magnetic field will be considered in a future version of SpectrAx. The inference of the magnetic field using synchrotron polarization measurements will provide better constraints on the magnetic field. Using the Faraday rotation $RM$ measurements, the magnetic field profile along the line of sight can also be constrained, which will be able to provide a 3-dimensional inference of the magnetic field direction and magnitude within the cluster. This can be included in an upgraded version of the SpectrAx code.
@@ 4.3 Framework for inference of the cluster electron density using X-rays
The estimation of 6 parameters ($n0$, $r_{c1}$, $\beta_{1}$, $T_{e}$, $\Gamma$ and $A$) using X-ray emission is shown is Fig. 6. The average time for the runs is $\approx 35$ core-hours per 1000 pixels. The electron density and temperature profiles involve a large number of parameters and some of them portray degeneracies. The constraints depend on the strength of the signal as well as the number of pixels that the cluster occupies. The 95% confidence region for the inference of electron density for the cluster is shown in Fig. 7.
@@
We have assumed that the electron density varies over the angular scale, which is resolvable with the beam resolution of eROSITA. We have considered a radial profile, but in reality the clusters will have a different geometry and the profiles will not be radial. This would require complex models with 3-dimensional variation of electron density. The turbulence in electron density will be considered in a future version of SpectrAx. SpectrAx works efficiently in estimating a large number of parameters, given computer resources and resolutions are available. The constrained parameters will be marginalized along with the magnetic field parameters and redshift to obtain the ALP coupling constant under the hierarchical Bayesian framework in Sec. 4.6.
SpectrAx will be able to employ complex profile models in its code and estimate the parameters. If the profile is simple, accompanied with good-quality data, the estimation will be robust and efficient. If the model is quite complex with a large number of parameters, some of the parameters may portray degeneracies. These degeneracies in profile parameters represent barriers in the way of constraining ALPs, due to astrophysical uncertainties. These can be dealt with by using better and multiple independent cluster observations from various surveys. The three dimensional profile of electron density and magnetic field will be considered in a future version of SpectrAx.
4.4 ALP signal in CMB map